OpenStudy (anonymous):

Jodie bikes 4 km/h faster than Dennis. In the same time it takes Dennis to bike 36 km, Jodie can bike 48 km. How fats does each bicyclist travel? Dennis= ? km/h Jodie= ? km/h

OpenStudy (anonymous):

*fast

OpenStudy (dumbcow):

distance = rate*time Dennis: 36 = rt Jodie 48 = (r+4)t = rt +4t use substitution replace rt in 2nd equation with 36 48 = 36 +4t 4t = 12 t = 3

OpenStudy (dumbcow):

oops not done yet so r = 36/t = 36/3 = 12

OpenStudy (anonymous):

dennise =12km/h jodie =16km/h